OpenStudy (anonymous):

Solve. x2 + 8x + 25 = 0

jimthompson5910 (jim_thompson5910):

Use the quadratic formula to solve. x = (-b+-sqrt(b^2-4ac))/(2a) x = (-(8)+-sqrt((8)^2-4(1)(25)))/(2(1)) x = (-8+-sqrt(64-(100)))/(2) x = (-8+-sqrt(-36))/2 x = (-8+sqrt(-36))/2 or x = (-8-sqrt(-36))/2 x = (-8+6*i)/2 or x = (-8-6*i)/2 x = -4+3i or x = -4-3i So the solutions are x = -4+3i or x = -4-3i where i = sqrt(-1)

OpenStudy (anonymous):

you could try doing thrugh the trial and error method of solving quadratic equations x +3 l +3x x +5 l +5x x^2 +25 l + 8x

OpenStudy (anonymous):

so factors are (x+3)(x+5)

OpenStudy (precal):